# Ceresa cycles of $X_{0}(N)$

### v1

@@ 3.1. Torsion Subgroups of Modular Jacobians
Throughout this section $p\geq 5$ denotes a prime number. The proof of Theorem 1 focuses on proving that the shadow $D$ of some Hecke operator on $J_{0}(p)$ is non-torsion. The rational torsion subgroup of $J_{0}(p)$ is completely understood due to the classical work of Mazur and Ogg. In this section we give an overview of some well known results. Note that $X_{0}(p)$ has two cusps, both rational, which we denote by $c_{0}$ and $c_{\infty}$.
We first recall the following proposition due to Ogg [30].
Proposition 9 (Ogg).
@@
In the same work, Ogg conjectured that the linear equivalence of this divisor generates the entire rational torsion subgroup of $J_{0}(p)$. This was proved by Mazur using the theory of Eisenstein ideals [29, Page 142].
Theorem 10 (Mazur).
@@ 3.3. The Minimal Regular Models of $X_{0}(p)$ and $X_{0}^{+}(p)$
A key ingredient of the proof of Theorem 1 is reduction modulo $p$. We give a brief description of the special fibres of the minimal regular models of $X_{0}(p)$ and $X_{0}^{+}(p)$ over $\mathbb{F}_{p}$.
The reductions of $X_{0}\left(p\right)$ and $X_{0}^{+}\left(p\right)$ are smooth away from $p$. The special fibre of the minimal regular model of $X_{0}(N)$ at $p|N$, where $N$ is square-free and coprime to $6$, is described in [12] and [29, Appendix]. We recall this for prime level $N=p\geq 5$
@@ 4. Proof of Theorem 1
The proof of Theorem 1 reduces to proving that the scaled shadow of some carefully chosen Hecke operator defines a point of infinite order in the Jacobian. We consider various points, principally $\operatorname{Sh}(T_{2})$, but also $\operatorname{Sh}(T_{3})$ and $\operatorname{Sh}(w_{p})$ depending on the congruence class of $p$. We denote this point by $D$ and note the following trivial corollary of Theorems 10 and 11.
@@ 4.1. Reduction Arguments
and since $J$ is the Jacobian of a curve, we can determine the component group from the special fibre of the minimal regular model of the curve, see [5, Section 9.6, Theorem 1]. As the special fibre described in Theorem 17 has a single component we deduce that $\Phi$ is trivial, and hence $\mathcal{J}(\mathbb{F}_{p})=\mathcal{J}^{0}(\mathbb{F}_{p})$. Moreover, by [5, Section 9.5, Theorem 1], $\mathcal{J}^{0}(\mathbb{F}_{p})\cong\operatorname{Pic}^{0}(\mathcal{X}_{/\mathbb{F}_{p}})$ and $\mathrm{red}$ is simply the composition
@@ 4.4. Primes congruent to 11 modulo 12
When $p\equiv 3\mod{8}$, $h\left(-4p\right)=3h\left(-p\right)$, and repeating the above we get $p\geq 3754815$, which we reduce to $49451$, and $266$ primes to check. The list of exceptions includes the primes $\{11,23,47,59,71\}$, for which $X_{0}(p)$ is hyperelliptic. For all non-hyperelliptic exceptions we find a supersingular $j$-invariant defined over $\mathbb{F}_{p^{2}}\setminus\mathbb{F}_{p}$ and argue as in Proposition 20. These cases are checked in the file primes11AL.m, using that the $j$-invariants of points in $D_{p}$ and $D_{4p}$ reduce to the $\mathbb{F}_{p}$-rational supersingular $j$-invariants modulo $p$.
@@ 5.2. Explicit Calculations
For $N\in\{11^{2}$, $5^{3}$, $13^{2}\}$, we work with $D=\operatorname{Sh}(w_{N})$. We begin by computing a canonical model for the curve using a basis of the space of weight $2$ cusps forms for $\Gamma_{0}(N)$, see [18]. In particular, we choose a basis whose elements are invariant or anti-invariant under the action of $w_{N}$ and from this we are able easily construct $D$ explicitly. We show that modulo different primes, the divisor has different orders, and hence by the injectivity of torsion [22, Appendix], we conclude that the shadow is necessarily of infinite order. See the online repository for these explicit calculations.
@@ 5.3. Period Lattice Computations
For $N\in\{3^{5},7^{3},17^{2},19^{2},29^{2},31^{2},37^{2},41^{2}\}$, the genus of $X_{0}(N)$ is large and computing explicit divisors in Magma is extremely inefficient. Instead, we use the analytic structure of the curve and its Jacobian, and project our shadow point into a (complex) subvariety of dimension $1$, $2$ or $3$. We give a short overview of the analytic definition of the Jacobian $J_{0}(N)$ and its modular subvarieties, see [13, Chapter 6] or [8, Chapter 2] for comprehensive descriptions.
@@
see [22, Appendix]. We may therefore take $B(N)$ to be the lowest common multiple of $|A_{f}(\mathbb{F}_{l})|$, as $l$ ranges over a finite set of odd primes coprime to $N$. Moreover, the cardinality of $A_{f}(\mathbb{F}_{l})$ can be determined directly from the coefficients of $f$ as detailed in [1, Section 3.5].

### v2

@@ 3.1. Torsion Subgroups of Modular Jacobians
Throughout this section $p\geq 5$ denotes a prime number. The proof of Theorem 1 focuses on proving that the shadow $D$ of some Hecke operator on $J_{0}(p)$ is non-torsion. The rational torsion subgroup of $J_{0}(p)$ is completely understood due to the classical works of Mazur and Ogg. In this section we give an overview of some well known results. Note that $X_{0}(p)$ has two cusps, both rational, which we denote by $c_{0}$ and $c_{\infty}$.
We first recall the following result due to Ogg [34].
Theorem 6 (Ogg).
@@
In the same work, Ogg conjectured that the linear equivalence of this divisor generates the entire rational torsion subgroup of $J_{0}(p)$. This was proved by Mazur using the theory of Eisenstein ideals [33, Page 142].
Theorem 7 (Mazur, Theorem III.1.2).
@@ 3.3. The Minimal Regular Models of $X_{0}(p)$ and $X_{0}^{+}(p)$
A key ingredient of the proof of Theorem 1 is reduction modulo the only prime of bad reduction for $X_{0}(p)$ and $J_{0}(p)$, namely $p$. We give a brief description of the special fibres of the minimal regular models of $X_{0}(p)$ and $X_{0}^{+}(p)$ over $\mathbb{F}_{p}$.
The reductions of $X_{0}\left(p\right)$ and $X_{0}^{+}\left(p\right)$ are smooth away from $p$. The special fibre of the minimal regular model of $X_{0}(N)$ at $p|N$, where $N$ is square-free and coprime to $6$, is described in [14] and [33, Appendix]. We recall this for prime level $N=p\geq 5$
@@
4. Fixed Points of Hecke Operators
In this section we describe how to compute the shadow points corresponding to Hecke operators of modular curves. These form the backbone of the proofs of Theorems 1 and 2. More precisely, for $C=X_{0}(N)$ the endomorphism ring of the Jacobian contains the Hecke algebra generated (over $\mathbb{Z}$) by the induced Atkin-Lehener operators $w_{d}$ for all $d|N$ and coprime to $\frac{N}{d}$; and for all $\ell\nmid N$ the Hecke operator $T_{\ell}$. For the rest of the paper we write $\mathbb{T}_{N}$ for the Hecke algebra.
Notation.
(1) $\operatorname{Sh}(\phi)=(2g-2)F_{\phi}-\deg(F)K_{C}.$
The fixed points of Atkin-Lehner operators are well-know due to Ogg [35, Proposition 3].
Proposition 15.
The points in the support of $F_{w_{d}}$ are either:
  • cusps if $d=4$;
  • CM points with endomorphism ring containing $\mathbb{Z}[\sqrt{-d}]$
Moreover, all fixed points occur with multiplicity $1$.
${X_{0}(N\ell)}$${X_{0}(N\ell)}$${X_{0}(N)}$${X_{0}(N)}$$\scriptstyle{w_{\ell}}$
see [15, Chapter 5] for details. As a result of this, we see that in the sense of the previous section, $T_{\ell}=T_{\ell}^{\vee}$. Further, as is well known, the degrees for both are $\ell+1$. This simplifies $\operatorname{Sh}(T_{\ell})$ to $(2g-2)F_{T_{\ell}}-\deg(F_{T_{\ell}})K_{X_{0}(N)}-2T_{\ell}(K_{X_{0}(N)})+2(\ell+1)K_{X_{0}(N)}$.
Proposition 16.
The points in the support of $F_{T_{\ell}}$ are either:
  • Cusps;
Proof.
Proposition 17.
Suppose $P\in X_{0}(N\ell)$ has both projections to $X_{0}(N)$ equal, then the image of the tangent space of $P$ under the map from $X_{0}(N\ell)$ to $X_{0}(N)^{2}$ is not diagonal.
Proof.
Otherwise, $P$ corresponds to an elliptic curve. By the previous proposition, this elliptic curve has CM. Since the $j$-invariant is a local parameter for CM points on $X_{0}(N)$, it is enough to check that $j-j\circ\omega_{\ell}$ has simple zeros at CM points on $X_{0}(N\ell)$. As this function is a pullback from $X_{0}(\ell)$, we start by proving this for $X_{0}(\ell)$. For $\ell=2$, $j-j\circ\omega_{\ell}$ has two poles of order 2, and so is a degree 4 function. There are 3 CM orders with an element of norm 2, $\mathbb{Z}[\sqrt{-1}]$, $\mathbb{Z}[\sqrt{-2}]$ and $\mathbb{Z}[\frac{-1+\sqrt{-7}}{2}]$. The first two of these have a unique ideal of index 2, and the second has 2 choices, giving rise to a total of 4 points on $X_{0}(2)$ where this function vanishes. This forces all 4 points to be simple zeros. An identical counting argument applies for $\ell=3$.
For $\ell\geq 5$, we can reduce modulo $\ell$. On the regular model of $X_{0}(p)$ modulo $\ell$, the Atkin-Lehner involution acts by swapping the Frobenius and Verschiebung components, and maps a $j$-invariant on the Frobenius component to $\operatorname{Fr}(j)$, and similarly for the Verschiebung component.
In particular, the function $j-j\circ\omega_{\ell}$ vanishes on the points of each component with $\mathbb{F}_{\ell}$-rational $j$-invariant. As there are $l$ such $j$-invariants, there would be at least $2\ell$ zeroes if each appears on a separate component. However, there are rational supersingular $j$-invariants, which correspond to complex multiplication by $\mathbb{Z}[\sqrt{-p}]$ and $\mathbb{Z}\left[\frac{-1+\sqrt{-p}}{2}\right]$. These CM curves reduce in pairs to the rational supersingular $j$-invariants, and so $j-j\circ\omega_{\ell}$ has all its zeroes in pairs, and one pair for each $j$-invariant over $\mathbb{F}_{\ell}$. In particular, it has at least $2\ell$ distinct zeroes, and has degree $2\ell$, and so all zeroes must be simple. ∎
Proposition 18.
Proof.
Combining the above in the special case of $T_{2}$, gives the following
$F_{T_{2}}=D_{4}+D_{8}+2D_{7}+2D_{\infty}$
@@ 5. Proof of Theorem 1
The proof of Theorem 1 reduces to proving that the scaled shadow of some carefully chosen Hecke operator defines a point of infinite order in the Jacobian. We consider various points, principally $\operatorname{Sh}(T_{2})$, but also $\operatorname{Sh}(T_{3})$ and $\operatorname{Sh}(w_{p})$ depending on the congruence class of $p$. We denote this point by $D$ and note the following trivial corollary of Theorems 7 and 8.
@@ 5.1. Reduction Arguments
and since $J$ is the Jacobian of a curve, we can determine the component group from the special fibre of the minimal regular model of the curve, see [6, Section 9.6, Theorem 1]. As the special fibre described in Theorem 14 has a single component we deduce that $\Phi$ is trivial, and hence $\mathcal{J}(\mathbb{F}_{p})=\mathcal{J}^{0}(\mathbb{F}_{p})$. Moreover, by [6, Section 9.5, Theorem 1], $\mathcal{J}^{0}(\mathbb{F}_{p})\cong\operatorname{Pic}^{0}(\mathcal{X}_{/\mathbb{F}_{p}})$ and $\mathrm{red}$ is simply the composition
@@ 5.4. Primes congruent to 11 modulo 12
When $p\equiv 3\mod{8}$, $h\left(-4p\right)=3h\left(-p\right)$, and repeating the above we get $p\geq 3754815$, which we reduce to $49451$, and $266$ primes to check. The list of exceptions includes the primes $\{11,23,47,59,71\}$, for which $X_{0}(p)$ is hyperelliptic. For all non-hyperelliptic exceptions we find a supersingular $j$-invariant defined over $\mathbb{F}_{p^{2}}\setminus\mathbb{F}_{p}$ and argue as in Proposition 21. These cases are checked in the file primes11AL.m, using that the $j$-invariants of points in $D_{p}$ and $D_{4p}$ reduce to the $\mathbb{F}_{p}$-rational supersingular $j$-invariants modulo $p$. A suitable $j$-invariant cannot be found for the bi-elliptic $X_{0}(p)$, and so are checked individually using the $T_{2}$ shadow.
@@ 6.2. Explicit Calculations
For $N\in\{11^{2}$, $5^{3}$, $13^{2}\}$, we work with $D=\operatorname{Sh}(w_{N})$. We begin by computing a canonical model for the curve using a basis of the space of weight $2$ cusps forms for $\Gamma_{0}(N)$, see [20]. In particular, we choose a basis whose elements are invariant or anti-invariant under the action of $w_{N}$ and from this we are able easily construct $D$ explicitly. We show that modulo different primes, the divisor has different orders, and hence by the injectivity of torsion [26, Appendix], we conclude that the shadow is necessarily of infinite order. See the online repository for these explicit calculations.
@@ 6.3. Period Lattice Computations
For $N\in\{3^{5},7^{3},17^{2},19^{2},29^{2},31^{2},37^{2},41^{2}\}$, the genus of $X_{0}(N)$ is large and computing explicit divisors in Magma is extremely inefficient. Instead, we use the analytic structure of the curve and its Jacobian, and project our shadow point into a (complex) subvariety of dimension $1$, $2$ or $3$. We give a short overview of the analytic definition of the Jacobian $J_{0}(N)$ and its modular subvarieties, see [15, Chapter 6] or [10, Chapter 2] for comprehensive descriptions.
@@
see [26, Appendix]. We may therefore take $B(N)$ to be the lowest common multiple of $|A_{f}(\mathbb{F}_{l})|$, as $l$ ranges over a finite set of odd primes coprime to $N$. Moreover, the cardinality of $A_{f}(\mathbb{F}_{l})$ can be determined directly from the coefficients of $f$ as detailed in [1, Section 3.5].
@@
7. Hecke Components of the Modified Diagonal
We specialise the results of Section 2.3 to the Hecke algebra $\mathbb{T}_{N}$ acting on $X=X_{0}(N)$. Throughout this section we write $\Gamma^{3}\in CH_{1}(X_{0}(N)^{3})$ for the modified diagonal cycle, with base divisor $\frac{K_{C}}{2g-2}$.
Recall that $\mathbb{T}^{\otimes 3}$ naturally acts on $CH_{1}(X_{0}(N)^{3})$, and that for any $T\in\mathbb{T}_{N}$, we have an associated shadow point
$\operatorname{Sh}(T)=\operatorname{Sh}(\Gamma^{3},T)\in J_{0}(N)(\mathbb{Q})$.
As simple consequence of Proposition 5 is the following.
Corollary 28.
For any Hecke operators $\phi_{1},\phi_{2},\phi_{3},T\in\mathbb{T}_{N}$,
$\operatorname{Sh}((\phi_{1},\phi_{2},\phi_{3})_{*}\Gamma^{3},T)=\phi_{3}\cdot\operatorname{Sh}(\Gamma^{3},\phi_{2}T\phi_{1}).$
By diagonalising the Hecke action on $CH_{1}(X_{0}(N)^{3})$, we decompose the modified diagonal cycle into Hecke isotypic components
$\Gamma^{3}=\displaystyle\sum_{f,g,h}\Gamma_{fgh}^{3}$
where the sum is taken over weight $2$ cusp forms of level $\Gamma_{0}(N)$. As $\operatorname{Sh}$ is linear, we obtain a corresponding decomposition of shadow points
$\operatorname{Sh}(\Gamma^{3},T)=\displaystyle\sum_{f,g,h}\operatorname{Sh}(\Gamma^{3}_{fgh},T)$
for any $T\in\mathbb{T}_{N}$. A natural question is to ask which of these components is non-trivial.
Proposition 29.
Let $f,g,h\in\mathcal{S}_{2}(\Gamma_{0}(N)$. If $f\neq g$, then $\operatorname{Sh}(\Gamma^{3}_{fgh},T)=0$ for all $T\in\mathbb{T}$.
Proof.
For any prime $q$, not dividing $N$, we have
$\operatorname{Sh}((T_{q},1,1)_{*}\Gamma^{3}_{fgh},T)=T_{q}\operatorname{Sh}(\Gamma^{3}_{fgh},T)=a_{f}(q)\operatorname{Sh}(\Gamma^{3}_{fgh},T).$
Equivalently, by definition
$\operatorname{Sh}((T_{q},1,1)_{*}\Gamma^{3}_{fgh},T)=\operatorname{Sh}(\Gamma^{3}_{fgh},T_{q}T)=\operatorname{Sh}((1,T_{q},1)_{*}\Gamma^{3}_{fgh},T)=a_{g}(q)\operatorname{Sh}(\Gamma^{3}_{fgh},T).$
Thus if $\operatorname{Sh}(\Gamma^{3}_{fgh},T)\neq 0$, $a_{f}(q)=a_{g}(q)$ for all $q$, and so $f=g$. ∎
We note that this does not force the vanishing of $\Gamma^{3}_{fgh}$ for $f\neq g$, simply that this does not contribute to shadows.
A special case occurs in the case of an Atkin-Lehner operator $w_{d}$ and the diagonal $T=\Delta$ on $X_{0}(N)^{2}$, i.e the trivial element in $\mathbb{T}_{N}$. As $w_{d}\otimes w_{d}\otimes w_{d}$ acts trivially on $\Gamma^{3}$, we decompose $\Gamma^{3}$ according to the sign with which $w_{d}$ acts on $\Gamma^{3}$
$\Gamma^{3}=\Gamma^{3}_{+++}+\Gamma^{3}_{+--}+\Gamma^{3}_{-+-}+\Gamma^{3}_{--+}$.
Proposition 30.
If $\operatorname{Sh}(\Gamma^{3},w_{d})\neq 0$, then $\operatorname{Sh}(\Gamma^{3}_{+++},\Delta)\neq 0$ and $\operatorname{Sh}(\Gamma^{3}_{--+},\Delta)\neq 0$. Moreover, all 4 Atkin-Lehner components of $\Gamma^{3}$ are non-zero. Conversely, if $\operatorname{Sh}(\Gamma^{3},w_{d})=0$ then $\operatorname{Sh}(\Gamma^{3}_{+++},\Delta)=0$ and $\operatorname{Sh}(\Gamma^{3}_{--+},\Delta)=0$.
Proof.
Using the above decomposition we obtain
$\operatorname{Sh}((w_{d}\otimes 1\otimes 1)_{*}\Gamma^{3},\Delta)=\operatorname{Sh}(\Gamma^{3}_{+++},\Delta)+\operatorname{Sh}(\Gamma^{3}_{+--},\Delta)-\operatorname{Sh}(\Gamma^{3}_{-+-},\Delta)-\operatorname{Sh}(\Gamma^{3}_{--+},\Delta);$
$\operatorname{Sh}((1\otimes w_{d}\otimes 1)_{*}\Gamma^{3},\Delta)=\operatorname{Sh}(\Gamma^{3}_{+++},\Delta)-\operatorname{Sh}(\Gamma^{3}_{+--},\Delta)+\operatorname{Sh}(\Gamma^{3}_{-+-},\Delta)-\operatorname{Sh}(\Gamma^{3}_{--+},\Delta).$
From Corollary 28, we deduce that both of the above divisors are $\operatorname{Sh}(\Gamma^{3},w_{d})$ and thus
$\operatorname{Sh}(\Gamma^{3},w_{d})=\operatorname{Sh}(\Gamma^{3}_{+++},\Delta)-\operatorname{Sh}(\Gamma^{3}_{--+},\Delta)$
Moreover, we have
$\operatorname{Sh}((1\otimes 1\otimes w_{d})_{*}\Gamma^{3},\Delta)=\operatorname{Sh}(\Gamma^{3}_{+++},\Delta)-\operatorname{Sh}(\Gamma^{3}_{+--},\Delta)-\operatorname{Sh}(\Gamma^{3}_{-+-},\Delta)+\operatorname{Sh}(\Gamma^{3}_{--+},\Delta);$
$\operatorname{Sh}(\Gamma^{3},\Delta)=\operatorname{Sh}(\Gamma^{3}_{+++},\Delta)+\operatorname{Sh}(\Gamma^{3}_{+--},\Delta)+\operatorname{Sh}(\Gamma^{3}_{-+-},\Delta)+\operatorname{Sh}(\Gamma^{3}_{--+},\Delta);$
and since shadow points are invariant under $w_{d}$ and $\operatorname{Sh}(\Gamma^{3},\Delta)=0$,
$\operatorname{Sh}(\Gamma^{3}_{+++},\Delta)=-\operatorname{Sh}(\Gamma^{3}_{--+},\Delta);$
and the claim follows for $\Gamma^{3}_{+++}$ and $\Gamma^{3}_{--+}$. The claim for the other two components follows by the invariance of $\Gamma^{3}$ under permutation of the factors. ∎
One motivation for studying this decomposition is the following conjecture of Gross and Kudla. For any three weight $2$ cusps forms $f,g,h$ of level $\Gamma_{0}(N)$, with $N$ squarefree, consider $F=f\times g\times h$ on $\mathbb{H}^{3}$. The associated $L$-function has been classically studied by Garrett [21] and Piatetski-Shapiro and Rallis [37], in particular it is centered at $s=2$. Associated to this setting, there is a canonically defined Shimura curve $X$, and as before we write $\Gamma^{3}$ for the modified diagonal cycle attched to $X$. Gross and Kudla [23] conjecture the following.
Conjecture.
With notation as above
$L^{\prime}(F,2)=\Omega(F)\langle\Gamma^{3}_{F},\Gamma^{3}_{F}\rangle_{BB}$
where $\Omega(F)$ is an explicit,non-zero constant, $\Gamma^{3}_{F}$ is the $F$-isotypic component of $\Gamma^{3}$ and $\langle,\rangle_{BB}$ is the Bloch-Beilinson height pairing.
There has been some recent progress towards proving this conjecture, see [43]. We provide some examples to support this conjecture where we prove the non-vanishing of $\Gamma^{3}_{F}$ and $L^{\prime}(F,2)\neq 0$. For instance, suppose that $X=X_{0}(p)$ is bielliptic with non-vanishing Ceresa cycle, that is, $p\in\{43,53,61,79,83,89,101,131\}$ (see [2]). For all such levels there is a unique $f_{p}\in\mathcal{S}_{2}(p)$ with Atkin-Lehner sign $1$ and in the following two results we exploit this to deduce that $\operatorname{Sh}(\Gamma^{3}_{ggf_{p}},\Delta)\neq 0$ for all $g\in\mathcal{S}_{2}(p)$, $g\neq f_{p}$.
Proposition 31.
Let $X=X_{0}(p)$ with $p\in\{43,53,61,79,83,89,101,131\}$. Then $\operatorname{Sh}(\Gamma^{3},w_{p})=0$. In particular, $\Gamma^{3}_{f_{p}f_{p}f_{p}}$ does not contribute to shadow points.
Proof.
By the previous proposition, it suffices to show $\operatorname{Sh}(\Gamma^{3},w_{p})=0$. This follows from our definition of $\operatorname{Sh}(\Gamma^{3},w_{p})$ and the fact that the canonical divisor is the sum of the fixed points of $w_{p}$, since the curve is bi-elliptic. ∎
In fact, $\Gamma^{3}_{f_{p}f_{p}f_{p}}=0$ as it is the pullback of the diagonal cycle on the elliptic curve, but for our purposes the statement of the proposition suffices.
Proposition 32.
Let $X=X_{0}(p)$ and $f_{p}$ be as above. Then $\operatorname{Sh}(\Gamma^{3}_{ggf_{p}},\Delta)\neq 0$ for all cusps forms $g$ with Atkin-Lehner sign $-1$.
Proof.
By the previous result, $\operatorname{Sh}(\Gamma^{3},\Delta)=0$ and thus
$\sum_{g}\operatorname{Sh}(\Gamma^{3}_{ggf_{p}},\Delta)=0$
where the sum is taken over a basis for the space of weight 2 cusps forms with Atkin-Lehner sign $-1$. Moreover, for all primes $l\neq p$
$\operatorname{Sh}(\Gamma^{3},T_{l})=\displaystyle\sum_{g}a_{g,l}\operatorname{Sh}(\Gamma_{ggf_{p}}^{3},\Delta)$
where $a_{g,l}$ is the eigenvalue of $T_{l}$ acting on $g$. Let $d$ be the dimension of the $-1$ Atkin-Lehner eigenspace. Taking $d-1$ different primes $l$, with at least one $\operatorname{Sh}(\Gamma^{3},T_{l})\neq 0$ (such a prime exists from our proof of Theorem 1), gives a system of $d$ equations in $d$ variables. The corresponding matrix (whose entries are the eigenvalues of the chosen Hecke operators) is invertible, since each column corresponds to a different embedding of the number field over which the cusp forms are defined. It follows by elementary linear algebra that $\operatorname{Sh}(\Gamma_{ggf_{p}}^{3},\Delta)\neq 0$ for all $g$.
We now turn to the $L$-function side of the conjecture.
Proposition 33.
Let $p\in\{43,53,61,79,83,89,101,131\}$. For all $g\in\mathcal{S}_{2}(p)$ with Atkin-Lehner sign $-1$, $L^{\prime}(g\otimes g\otimes f_{p},2)\neq 0$.
Proof.
We make use of the factorisation $L(g\otimes g\otimes f_{p})=L(f_{p}\otimes\mathrm{Sym}^{2}g,s)L(f_{p},s-1)$ to reduce calculating $L^{\prime}(g\otimes g\otimes f_{p},2)$ to $L^{\prime}(f_{p},1)$ and $L(f_{p}\otimes\mathrm{Sym}^{2}g,2)$. These can be computed using the computer algebra package $\mathtt{Magma}$, after correcting the local $L$-factor at $p$. ∎
In theory, such calculations can be carried out for any $X_{0}(N)$.
Example.
Let $X=X_{0}(67)$. The $+$ space is spanned by two conjugate cusp forms $h_{1},h_{2}$ defined over $\mathbb{Q}(\sqrt{5})$; whilst the -- space is spanned by a rational cusp form $f$ and a pair of Galois conjugate forms $g_{1},g_{2}$ defined over $\mathbb{Q}(\sqrt{5})$.
We compute models of $X_{0}(67)$ and $X_{0}^{+}(67)$, along with an isomorphism $J_{0}^{+}(67)(\mathbb{Q})\cong\mathbb{Z}^{2}$. Shadow points can be determined explicitly using our models and various facts about quadratic points proved in [7]. We push-forward the divisors onto $J_{0}^{+}(67)(\mathbb{Q})$ to find:
$3\operatorname{Sh}(T_{2})=24(1,3),3\operatorname{Sh}(T_{3})=24(-10,3),3\operatorname{Sh}(T_{7})=24(-1,-7),3\operatorname{Sh}(w_{67})=48(1,-1),$
with respect to a fixed basis of $J_{0}^{+}(67)\cong\mathbb{Z}^{2}$. Along with $\operatorname{Sh}(\Delta)=0$, and the images of these 5 points under $T_{2}$, we obtain a system of 10 equations in the 10 variables $a_{s,t}=\operatorname{Sh}(\Gamma^{3}_{sst},\Delta)$ where $s\in\{h_{1},h_{2},f,g_{1},g_{2}\}$ and $t\in\{h_{1},h_{2}\}$. Explicitly solving this system, we find that $\operatorname{Sh}(\Gamma^{3}_{h_{1}h_{1}h_{1}},\Delta)=\operatorname{Sh}(\Gamma^{3}_{h_{2}h_{2}h_{2}},\Delta)=0$, which is consistent with the vanishing of the respective triple product $L$-function. All other Hecke components are non-zero, as are the corresponding derivatives of the associated $L$-function.